what is the answer to 3(m+1)+2(m+2)=52

Many Thanks
1
Expert's answer
2012-10-18T07:25:02-0400
3(m+1)+2(m+2)=52
Open the braces:
3m+3+2m+4=52
Gather terms with m in the left part, others in the right:
5m=52-7=45
m=9